OpenStudy (anonymous):

What is the cost in cents of one baseball if baseballs are priced at d dollars per dozen?

OpenStudy (anonymous):

12 baseballs = d dollars 12/12 baseball = d/12 dollars 1 baseball = d/12 dollars (d/12) * 100 = 100d/12 = 25d/3 = (25/3)d (25/3)d cents for 1 baseball

OpenStudy (anonymous):

where did u get 100 from?

OpenStudy (anonymous):

oh never mind i just figured out that its because there are 100 cents in a dollar